Home‎ > ‎Reports‎ > ‎Crew 108‎ > ‎

day10


Crew #108 Log Book for December 12, 2011

Photos of the day





Commander's Check-In Report

Date and time:

12 December 2011

Written by :

Charlotte Poupon, Crew 108, Commander

Crew Physical Status:

All ok

Time Departed/Returned from EVA:

EVA # 17 / 10:35 to 12:05
EVA # 18 / 10:35 to 13:00

Brief Narrative of Field Mission Results:

Not applicable

EVA Data/Interpretations:

We went for a pedestrian EVA to south to take more pictures of the neighborhooddue to the snow coming.

Engineering/Hab Maintenance

By this afternoon, engineers got in touch with Don to have our water tank refilled. Which was done by this evening. Thank you Don !

Report Transmission Schedule:

All reports by 22:00.

Plans for Tomorrow:

Hum... I suspect my buddy will go for a snow spaceman... And I'll be happy to help !

Inventory:

Not applicable

Support Requested:

We asked for stuff to build an Alternative Breathing System during EVA. See previous mail to Mission Support. Waiting for your answer.

Miscellaneous:

Commander's Report

Date and time: 

12 December 2011

Written by: 

Charlotte Poupon, Crew 108, Commander

We made a pedestrian EVA around the Hab, South Direction. During this
EVA we performed a test of a dysfunction in a breathing gear. This exercise was previously documented.

We are suggesting to think about this issue and offered to make a mock-up of a new Alternative Breathing System (ABS) that could be used in Sim First Aid.

Waiting a comment from Mission Support

EVA 17 Report

Date and time: 

12 December 2011

Written by: 

Ashley Dale, Crew 108, Engineer

Crew Members: Michael LeClair (Geologist), Usha Lingappa (Astrobiologist), Ashley Dale (Engineer), Charlotte Poupon (Station Leader), Mike Lotto (Engineer)

Time: 09:35 – 12:15

- Depressurization: 09:35 – 09:40

- “Failure of a respirator” emergency exercise: 10h12 – 10h17

- Separation of group (Group A: Poupon, Lingappa, Dale / Group B: Lotto, LeClair): 11h56

- Group A walk back to Hab: 11:56 – 12:08

- Repressurization for Group A: 12:08 – 12:13

- Dale becomes Hab Support for Group B: 12:15

- See EVA Report 18 for Group B part of excursion

Site Location: 

Transit Mode: 

Foot

Objectives: 

Explore region south of Hab for biological and geophysical interest, photo-opp for entire crew. Carry out “Failure of a respirator” simulation and record time taken to supply air from another crew member.

Results: 

No regions of biological interest encountered, geophysical interest experience and separation of EVA (see EVA 18 Report for details). “Failure of a respirator” simulation carried out successfully – took a little less than 2 minutes from “failure” to “resupply”. Great photos taken.

Lessons Learned: 

An “Alternative Breathing System” concept was conceived by Poupon, and proposed for prototyping to Mission Support. Failure of respiratory systems is a very real obstacle when on EVAs, and such a concept could quite easily be explored further at MDRS.

EVA 18 Report

Date and time: 

12 December 2011

Written by: 

Michael LeClair, Crew 108, XO

Crew Members: Michael LeClair (Geologist), Mike Lotto (Engineer)

Hab Support: Ashley Dale (Engineer)

Time: 09:35 – 12:57

- Depressurization: 09:35 – 09:40

- Separation of group (Group A: Poupon, Lingappa, Dale / Group B: Lotto, LeClair): 11h56

- Group A walk back to Hab: 11h56 – 12h08

- Repressurization for Group A: 12h08 – 12h13

- Dale becomes Hab Support: 12h15

- Group B walk back to Hab: 12h33 – 12h52

- Repressurization for Group B: 12h52 – 12h57

Site Location: The ‘Motherload’ waypoint -as documented by a previous Crew- was located where hematite concretions in highly chloriticly-altered stone, possibly sandstone, were evident.

Multiple obvious detrital sandstone pieces with abundant 0.25 to 0.5 cm hematitic concretions identified starting at base of ridge and moving up the rectilinear slope. These finds are directly below and down-slope from ridge line samples collected during EVA 12.

EVA continued south, climbing eastern-side cut of Kissing Camel Ridge to peak, close to Piece of Cake. Conglomerate capstone here closely resembles those of the rest of the ridge opposite the road cut to the east and continuing NE, terminating at Phobos Peak. This sinuous ridge feature is interpreted as an inverted paleo-channel.

Transit Mode: 

Foot

Objectives: 

Explore region down-slope from EVA 12 concretion find.

Results: 

Success

Lessons Learned: 

Field Geology is an iterative and incremental process and should not be rushed.

Engineering Report

Date and time:
 
12 December 2011, 18:45

Written by:

Mike Lotto, Crew 108, Engineer

Generator/Electricity:

Did you have any electricity outage? (Yes/No): No

Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0.375
Propane Tank (Full/0.75/0.5/0.25/Empty): 0.75
Gasoline Tank (Full/0.75/0.5/0.25/Empty): Full
Notes/Comments:

Oil:

ATV Oil Quantity (# of Quarts in storage): 1.25
Car Oil Quantity (# of Quarts in storage): 1.25

Water Status:

Outside Potable Water Tank Level (inches from bottom): 6
Trailer Potable Water Tank Level (inches from bottom): Full
Hab Potable Water Tank Level (inches from bottom): 10
Potable Water Meter Reading: Unknown
In to GreenHab Meter Reading: Offline
Notes/Comments: The trailer tank was filled today. The outside potable water tank will be filled tomorrow morning.

GreenHab:

Condition of plants in Tank 1: Offline
Condition of plants in Tank 2: Offline
Condition of plants in Tank 3: Offline
Greenhab Notes/Comments: GreenHab is offline this season.

Observatory:

Telescope Used? (Yes/No): No
Observatory Notes/Comments: We've completed the on-line training class, and we request the keys to the observatory.

Transportation:

Hab Car (NOT RENTAL)
Used (yes/no): No
Oil Checked (yes/no): No
Tire Status: Good
Odometer at end of day: 125464 miles
Notes/Comments on Hab Car:

ATV 1 (Opportunity):
Used: No
Oil Checked: No
Fuel Consumed: 0 gallons
Tire Status: Good
No. of hours ridden since last oil change (HH:MM): 03:07
Comments:

ATV 2 (Spirit):
Used: No
Oil Checked: No
Fuel Consumed: 0 gallons
Tire Status: Good
No. of hours ridden since last oil change (HH:MM): 01:11
Comments:

ATV 3 (Viking I):
Used: No
Oil Checked: No
Fuel Consumed: 0 gallons
Tire Status: Good
No. of hours ridden since last oil change (HH:MM): 03:07
Comments:

Heating and Ventilation:

Thermostat setting upstairs (Farenheit): 67
Actual temperature upstairs (Farenheit): 68

Thermostat setting downstairs (Farenheit): 55
Actual temperature downstairs (Farenheit): 45

Computers/Networking Infrastructure:

Notes/comments on internet/computers?: None

General Engineering Notes/Comments: None

Journalist Report

Date and Time: 

12 December 2011, 18:26

Written by: 

Usha Lingappa, Crew 108, Biologist

The first part of the day was dominated by 3 epic chess games, all of which were won by Mr. Ashley Dale. In mid-morning, we all left on an EVA to explore the area directly south of the Hab by foot. Dale, Poupon, and I returned round noon, and Mike and Mike returned about an hour later. 

After lunch, we all spent the afternoon in the Hab, and were visited by DG with a re-fill for the trailer tank. I have been battling with the incubator in the lab, trying to get it to hold a temperature just below its lowest setting. It is currently snowing, and it is forecasted to continue snowing tomorrow and the next day. Apart from mourning the loss of clear skies, we haven't let the weather dampen our spirits. The landscape outside the window looks more like the
Moon than Mars.